HELIUM NANODROPLET ISOLATION SPECTROSCOPY AND AB INITIO CALCULATIONS OF HO_{3}- (O_{2})_{N} CLUSTERS (N=0-4)

\underline{T. LIANG}, P. RASTON, and G. E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602, USA.

The HOOO hydridotrioxygen radical and its deuterated analog (DOOO) have been isolated in helium nanodroplets following the in-situ association reaction between OH and O_{2}. The infrared spectrum in the 3500-3700 cm^{−1} region reveals bands that are assigned to the ν_{1} (OH stretch) fundamental and ν_{1} + ν_{6} (OH stretch plus torsion) combination band of the trans-HOOO isomer. The helium droplet spectrum is assigned on the basis of a detailed comparison to the infrared spectrum of HOOO produced in the gas phase [E. L. Derro, T. D. Sechler, C. Murray, and M. I. Lester, J. Chem. Phys. 128, 244313 (2008)]. Despite the characteristic low temperature and rapid cooling of helium nanodroplets, there is no evidence for the formation of a weakly bound OH-O_{2} van der Waals complex, which implies the absence of a kinetically significant barrier in the entrance channel of the reaction. There is also no spectroscopic evidence for the formation of cis-HOOO, which is predicted by theory to be nearly isoenergetic to the trans isomer. Stark spectroscopy of the trans-HOOO species provides vibrationally averaged dipole moment components that qualitatively disagree with predictions obtained from CCSD(T) computations at the equilibrium, planar geometry, indicating a floppy complex undergoing largeamplitude motion about the torsional coordinate. Under conditions that favor the introduction of multiple O_{2} molecules to the droplets, bands associated with larger H/DOOO-(O_{2})_{n} clusters are observed shifted 1-10 cm^{−1} to the red of the trans-H/DOOO ν_{1} bands. Detailed ab initio calculations are carried out for multiple isomers of cis- and trans-HO_{3}-O_{2}, corresponding to either hydrogen or oxygen bonded van der Waals complexes. Comparisons to theory suggest that the structure of the HO_{3}-O_{2} complex formed in helium droplets is a hydrogen-bonded ^{4}A′ species consisting of a trans-HO_{3} core. The computed binding energy of the complex is approximately 240 cm^{−1}. Despite the weak interaction between trans-HO_{3} and O_{2}, non-additive red shifts of the OH stretch frequency are observed upon successive solvation by O_{2} to form the larger clusters with n>1.